OpenStudy (anonymous):

I'm having so trouble with percentages. Say the question is 75 is 20% of what number. Can you explain it?

OpenStudy (texaschic101):

" is" means equals and "of " means multiply 75 is 20% of what number 75 = .20x (turn your decimal to a percent) now you divide by .20 75/.20 = x 375 = x check... 75 = .20(375) 75 = 75 (correct)

OpenStudy (anonymous):

Thank you that helped alot!

OpenStudy (texaschic101):

no problem.....just remember, in problems like this " is " means equals and " of " means multiply

OpenStudy (anonymous):

percent meaning the per 'cent' ie per hundred. 100% percent is 100 out of cent ie out of 100 which is full 50 percent is 50 out of cent ie out of 100 which is half. generally when we say something lets say x is 50% of another thing lets say y, then we mean that if another thing ie y is a hundred then something which is x would be 50 . if y is other than hundred then x is y*(percentage)/100

OpenStudy (anonymous):

Alright rhanks

hero (hero):

Just remember this formula: \[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\] Then insert the proper numbers: \[\frac{75}{x} = \frac{20}{100}\]